OpenStudy (anonymous):

complete the solution of the equation. find the value of y when x equals -14. -6(-14)+6y=42 /\ 84+64=42 explain.

OpenStudy (anonymous):

@((*-*))

OpenStudy (triciaal):

i think you have an error in what you copied. -6x + 6y = 42 when x = -14 -6(-14) + 6Y = 42 6Y = 42 -84 = -42 Y = -42/6 = -7

OpenStudy (anonymous):

yeah something like that

OpenStudy (triciaal):

84+64=42 you wrote 64 instead of 6y

OpenStudy (anonymous):

oh i did wow lol well thanks for the help!

OpenStudy (anonymous):

ill be sure to look it over a few times to understand it

OpenStudy (triciaal):

please let me know if you have any questions about my post I will go over it with you.

OpenStudy (anonymous):

okay :)